# Caging in Time: A Framework for Robust Object Manipulation under Uncertainties and Limited Robot Perception

### v3

@@ 4.2 Caging in Time
Unlike the traditional caging introduced in Sec. 3.1, our Caging in Time framework only requires one robot to manipulate the object. The cage is formed over time by switching the single robot to a different configuration and interacting with the object differently at each time step. Fig. 6 shows an example where a robot pusher can push an object from various possible initial locations relative to the object (marked by the grey bars in the right figure). At each time step, by predicting the object’s possible motion via the propagation function $\Pi$ defined in Sec. 4.1, the robot needs to figure out which candidate push to select (e.g., the green bar in the right figure of Fig. 6), to prevent the object from escaping. As such, the object can always be confined to a bounded region as the robot pusher switches to different positions over time and pushes the object in different directions.
Specifically, for every time step $t$, we want to confine the object’s state to be always inside a region $\mathcal{S}_{cage}^{t}\subset\mathcal{S}_{obj}$ in its state space, and we term this region as the “cage region”. The “cage region” $\mathcal{S}_{cage}^{t}$ does not have to be stationary; it can deform and displace depending on the manipulation task it entails. Moreover, a set of candidate robot actions $\mathcal{U}$ needs to be pre-defined based on the context of the manipulation task. For example, in the planar pushing manipulation scenario shown in Fig. 6, $\mathcal{U}$ can be a set of robot-pushing actions defined by different starting positions and pushing directions of the robot pusher. Such robot actions will affect the possible motion of the object. For example, while the robot is interacting with the object through contact, the object will not be able to move in the directions that penetrate the robot. In other words, for a specific object state $\mathbf{q}_{t}\in\mathcal{S}_{obj}$, the allowed object motions $\mathcal{V}_{\mathbf{q}_{t}}$ will be influenced by the robot action $u\in\mathcal{U}$. We use a function $U$ to represent such influence by
@@
The process of determining whether Caging in Time is achieved is detailed in Alg. 1. When the Caging in Time condition is met, the object is guaranteed to be caged inside a cage region $\mathcal{C}_{cage}^{t}$ through a certain sequence of robot actions in an open-loop manner. If we apply this sequence of robot actions, the object will never escape even without any sensing feedback. As aforementioned, the cage region $\mathcal{S}_{cage}^{t}$ does not have to be static, it is allowed to change over time. For example, if $\mathcal{S}_{cage}^{t}$ is moved to follow a certain trajectory in the state space of the object over time, the object being caged in time will be able to follow the same trajectory of $\mathcal{S}_{cage}^{t}$.
@@
Input: The time-varying cage region $\mathcal{S}_{cage}^{t}\subset\mathcal{S}_{obj}$, a sequence of robot actions $\{u_{t}\in\mathcal{U}\}$
Output: The object is caged in time or not
@@ 5 Quasi-static Tasks
In this section, we instantiate the Caging in Time theory on a quasi-static planar pushing problem. To facilitate the instantiation, we also develop relevant tools for propagating the PSS of the object in a 2D space and generating open-loop robot pushing actions to cage the object. With the Caging in Time framework applied to the planar pushing problem, we can enable a single robot to push an object of unknown shape to follow certain trajectories in an open-loop manner, without requiring sensory feedback, exact geometric information, or any other physical properties of the object. This is an instantiation of Caging in Time in a geometry-based state space.
@@ 5.1 Problem Statement
While the configuration space of a quasi-static 2D object is normally $SE(2)$, requiring both the position and orientation of the object, our use of a bounding circle allows us to simplify the object’s state space to its position only: $\mathcal{S}_{obj}\subset\mathbb{R}^{2}$. As such, the state of the object becomes $\mathbf{q}_{t}=(x_{t},y_{t})\in\mathcal{S}_{obj}$. Therefore, both the PSS $\mathcal{Q}_{t}$ and the cage $\mathcal{S}_{cage}^{t}$ can be conceptualized as regions within this 2D plane. In this specific planar pushing task, we further define the Potential Object Area (POA), $\mathcal{P}_{t}\subset\mathbb{R}^{2}$, to be the set consisting of all the points in the workspace that are possibly occupied by the object’s geometric shape:
@@
The robot action is characterized by a scalar angle $\theta$, as shown in the left of Fig. 7. The robot will first place the center of the line pusher at a starting position determined by $\theta$. This starting position is always on the boundary of the circle $\mathcal{S}_{\mathcal{A}}^{t}$ and can be calculated by $(x^{c}_{t}+R\cos{\theta},y^{c}_{t}+R\sin{\theta})$. The line pusher will be oriented to be tangent to the boundary circle of $\mathcal{S}_{\mathcal{A}}^{t}$. Then, the robot will move the pusher in the direction towards the center of $\mathcal{S}_{\mathcal{A}}^{t}$ with a fixed distance $d_{push}$ to push the object. We assume there are $K$ candidate robot actions, i.e., $\mathcal{U}=\{\theta_{k}\}_{k=1}^{K}$. The value of each action is obtained through $\theta_{k}=2k\pi/K$ such that the starting positions of these $K$ candidate actions will evenly surround $\mathcal{S}_{\mathcal{A}}^{t}$.
As defined in Sec. 4, we require the object’s state to be always confined inside the circular cage region $\mathcal{S}_{cage}^{t}$, i.e., the PSS of the object $\mathcal{Q}_{t}\subset\mathcal{S}_{cage}^{t}$. Then, for the object to follow a certain trajectory $\mathcal{T}=\{(x^{d}_{1},x^{d}_{1}),\cdots,(x^{d}_{t},y^{d}_{t}),\cdots,(x^{d}_{T},y^{d}_{T})\}$ represented by a sequence of waypoints (i.e., desired positions) in the object’s state space, we just need to move the center of $\mathcal{S}_{cage}^{t}$, which is $(x^{c}_{t},y^{c}_{t})$, along the same trajectory. As the object is always confined inside the moving $\mathcal{S}_{cage}^{t}$ for every time step $t=1,\cdots,T$, the object’s position is guaranteed to follow the trajectory $\mathcal{T}$ with a positional error no greater than the cage size $R-r$.
@@ 5.2 Unknown Object Shape and Its Motion
For planar pushing, we represent the object’s motion by the displacement of the object’s position between adjacent time steps, which we denote as $\mathbf{v}_{t}=\Delta\mathbf{q}_{t}=(\Delta x_{t},\Delta y_{t})$. Then the propagation function $\pi$, initially defined in Sec. 4.1, for a single object’s configuration and motion as inputs becomes $\pi(\mathbf{q}_{t},\mathbf{v}_{t})=\pi(\mathbf{q}_{t},\Delta\mathbf{q}_{t})=\mathbf{q}_{t}+\Delta\mathbf{q}_{t}$, which is used to evaluate the propagated object’s configuration $q_{t+1}$ at the next step.
For a specific object’s configuration $\mathbf{q}_{t}$, the set of all possible motions $\mathcal{V}_{\mathbf{q}_{t}}$ is modeled by the interaction between the object and the robot’s line pusher $\mathcal{L}$ while the action $u_{t}$ is being executed, as represented by the function $U$ defined in Eq. (6). The line pusher approaches the object with a straight-line distance $d_{push}$ while $u_{t}$ is being executed. If the distance between the object’s configuration and the line pusher $\mathbf{d}(\mathbf{q}_{t},\mathcal{L})$ is greater than $r+d_{push}$, the object’s geometric shape (i.e., the bounding circle) will not collide with the line pusher. The object will not move at all, and thus $\mathcal{V}_{\mathbf{q}_{t}}=\left\{(0,0)\right\}$. However, if the object’s bounding circle collides with the line pusher, which indicates the object is likely displaced by the line pusher, we denote $d_{con}\leq d_{push}$ as the moving distance of the line pusher after contacting the bounding circle of the object. As shown in Fig. 8, we can derive a bound for this displacement using the friction theory and Peshkin Distance Peshkin and Sanderson (1988) as follows:
$\mathcal{V}_{\mathbf{q}_{t}}=U(\mathbf{q}_{t},u_{t})=$ (9)
if $\mathbf{d}(\mathbf{q}_{t},\mathcal{L})>r+d_{push}$ or $u_{t}$ is None:
@@
With this discretized representation, the PSS of the object is implemented by a finite set of pixels on $\mathcal{I}_{t}$. We also do a similar discretization for the possible motion set $\mathcal{V}_{\mathbf{q}_{t}}$. As such, to propagate the PSS on the discretized space, we just enumerately propagate every pixel in the PSS with every motion in the discretized finite set of $\mathcal{V}_{\mathbf{q}_{t}}$. As mentioned, $\mathcal{I}_{t}$ is always centered with the moving cage. So we need to deal with an offset of $(x^{c}_{t},y^{c}_{t})$ whenever we convert between a pixel location on $\mathcal{I}_{t}$ and its actual coordinates in the original continuous space. Alg. 3 outlines the discretized implementation of the PSS propagation.
@@ 5.3 Cage the Push in Time
Recall that in Sec. 5.1, each candidate action in $\mathcal{U}$ is represented by an angle $\theta_{k}$, which determines the starting pose of the line pusher. For each candidate action in $\mathcal{U}$ with an index $k=1,\cdots,K$, we will evaluate a heuristic score by $h^{k}=\lambda_{1}S_{out}^{k}+\lambda_{2}\left(d_{out}^{k}\right)^{2}$ where $\lambda_{1}$ and $\lambda_{2}$ are two weighting factors. As shown in the right of Fig. 7, $S_{out}^{k}$ is the POA area behind the line pusher when we place the pusher at a starting pose determined by the angle $\theta_{k}$, and $d_{out}^{k}$ is the furthest distance from behind the line pusher to the POA boundary. It is worth noting that calculating the heuristics $h^{k}$ for each action does not require the propagation of PSS, i.e., the PSS remains $\mathcal{Q}_{t}$. Intuitively, a large $h^{k}$ indicates that a large portion of the PSS will go outside the cage if the $k$-th action is not taken.
@@ 6 Dynamic Tasks
In this section, we extend our Caging in Time theory from quasi-static tasks to dynamic scenarios where we instantiate the same framework on a challenging dynamic ball balancing problem. Tools are developed for propagating the ball’s PSS of its position and velocity, as well as for generating open-loop robot actions to keep the ball balanced while the robot end-effector follows different trajectories without requiring any sensory feedback. This is an instantiation of Caging in Time in an energy-based state space.
@@ 6.1 Problem Statement
${\color[rgb]{0,0,0}E(\mathbf{q}_{t},\mathbf{g}_{\theta},\mathbf{a}_{p})}=E_{k}(\mathbf{\dot{x}}_{t})+E_{ve}(\mathbf{x}_{t})+{\color[rgb]{0,0,0}E_{p}(\mathbf{x}_{t},\mathbf{g}_{\theta},\mathbf{a}_{p})}$ (12)
@@
${\color[rgb]{0,0,0}E_{max}(\mathbf{g}_{\theta},\mathbf{a}_{p})}=E_{ve}(l)+{\color[rgb]{0,0,0}E_{p}(l,\mathbf{g}_{\theta},\mathbf{a}_{p})}$ (13)
@@ 6.2 Ball Dynamics and PSS Propagation
[…]
@@ 6.3 Cage the Ball in Time
To achieve Caging in Time for the dynamic ball balancing problem, we employ a control optimization strategy that combines Control Barrier Function (Ames et al., 2019) and Control Lyapunov Function (Blanchini, 1995; Anand et al., 2021) to keep the ball remain on the plate while following the desired trajectory.
In our case, Control Barrier Function (CBF) is utilized to implement the concept of “caging” in dynamic setting. Just as a physical cage constraining an object within a defined space, the CBF establishes an energy-based cage that can be characterized by the margin between the maximum energy attainable within the PSS and the predefined energy boundary:
@@
Figure 14: Experiment setup. Left: Robot and camera setup. The robot performs the experiment on a transparent surface (blue) with a pre-installed line pusher (green). We used cameras 2 and 3 below the surface to record (not to track) the trajectories of the objects through April-Tag. Camera 1 is used to record the experiment from an upper view. Top Right: The 3D-printed objects used in the experiments. 1. Square with two trimmed tips, 2. Pentagon, 3. Triangle with trimmed tips, 4. Octagon, 5. Square, 6. Heart, 7. Hexagon, 8. Ellipse, 9. Circle, and letter-shaped objects: R, I, C, E. Bottom Right: 3D printed plates, tennis ball and plastic fruits for dynamic tasks. The square plate is for ball balancing when dimension $n=2$, and the thinner one is for ball balancing and catching when dimension $n=1$, which has a guide rail to constrain the motion of the ball only on the required dimension. The plastic fruits are strawberry (YCB #12), plum (YCB #18), pear (YCB #16) and lemon (YCB #14).

### v4

@@ 4.2 Caging in Time
Unlike the traditional caging introduced in Sec. 3.1, our Caging in Time framework only requires one robot to manipulate the object. The cage is formed over time by switching the single robot to a different configuration and interacting with the object differently at each time step. Fig. 6 shows an example where a robot pusher can push an object from various possible initial locations relative to the object (marked by the grey bars on the right). At each time step, by predicting the object’s possible motion via the propagation function $\Pi$ defined in Sec. 4.1, the robot needs to figure out which candidate push to select (e.g., the green bar in the right figure of Fig. 6), to prevent the object from escaping. As such, the object can always be confined to a bounded region as the robot pusher switches to different positions over time and pushes the object in different directions.
Specifically, for every time step $t$, we want to confine the object’s state to be always inside a region $\mathcal{S}_{cage}^{t}\subset\mathcal{S}_{obj}$ in its state space, and we term this region the “cage region”. The “cage region” $\mathcal{S}_{cage}^{t}$ does not have to be stationary; it can deform and displace depending on the manipulation task it entails. Moreover, a set of candidate robot actions $\mathcal{U}$ needs to be predefined based on the context of the task. For example, in the planar pushing manipulation scenario shown in Fig. 6, $\mathcal{U}$ can be a set of robot-pushing actions defined by different starting positions and pushing directions of the robot pusher. Such robot actions will affect the possible motion of the object. For example, while the robot is interacting with the object through contact, the object will not be able to move in the directions that penetrate the robot. In other words, for a specific object state $\mathbf{q}_{t}\in\mathcal{S}_{obj}$, the allowed object motions $\mathcal{V}_{\mathbf{q}_{t}}$ will be influenced by the robot action $u\in\mathcal{U}$. We use a function $U$ to represent such influence by
@@
The process of determining whether Caging in Time is achieved is detailed in Alg. 1. When the Caging in Time condition is met, the object is guaranteed to be caged inside a cage region $\mathcal{S}_{cage}^{t}$ through a certain sequence of robot actions in an open-loop manner. If we apply this action sequence, the object will never escape even without any sensing feedback. As aforementioned, the cage region $\mathcal{S}_{cage}^{t}$ does not have to be static; it is allowed to change over time. For example, if $\mathcal{S}_{cage}^{t}$ is moved to follow a certain trajectory in the state space of the object over time, the object being caged in time will be able to follow the same trajectory of $\mathcal{S}_{cage}^{t}$.
@@
Input: The time-varying cage region $\mathcal{S}_{cage}^{t}\subset\mathcal{S}_{obj}$
, a sequence of robot actions $\{u_{t}\in\mathcal{U}\}$ Output: The object is caged in time or not
@@ 5 Quasi-static Tasks
In this section, we instantiate the Caging in Time theory on a quasi-static planar pushing problem. To facilitate the instantiation, we also develop relevant tools for propagating the PSS of the object in a 2D space and generating open-loop robot pushing actions to cage the object. With the Caging in Time framework applied to the planar pushing problem, we can enable a single robot to push an object of unknown shape to follow certain trajectories in an open-loop manner, without requiring sensory feedback, exact geometric information, or any other physical properties of the object. This is an instantiation of Caging in Time with a geometry-based cage.
@@ 5.1 Problem Statement
While the configuration space of a quasi-static 2D object is $SE(2)$, requiring both the position and orientation of the object, our use of a bounding circle allows us to simplify the object’s state space to its position only: $\mathcal{S}_{obj}\subset\mathbb{R}^{2}$. As such, the state of the object becomes $\mathbf{q}_{t}=(x_{t},y_{t})\in\mathcal{S}_{obj}$. Therefore, both the PSS $\mathcal{Q}_{t}$ and the cage $\mathcal{S}_{cage}^{t}$ can be conceptualized as regions within this 2D plane. In this specific planar pushing task, we further define the Potential Object Area (POA), $\mathcal{P}_{t}\subset\mathbb{R}^{2}$, to be the set consisting of all the points in the workspace that are possibly occupied by the object’s geometric shape:
@@
The robot action is characterized by a scalar angle $\theta$, as shown in the left of Fig. 7. The robot will first place the center of the line pusher at a starting position determined by $\theta$, which is always on the boundary of the circle $\mathcal{S}_{\mathcal{A}}^{t}$ and can be calculated by $(x^{c}_{t}+R\cos{\theta},y^{c}_{t}+R\sin{\theta})$. The line pusher will be oriented to be tangent to the boundary circle of $\mathcal{S}_{\mathcal{A}}^{t}$. Then, the robot will move the pusher in the direction towards the center of $\mathcal{S}_{\mathcal{A}}^{t}$ with a fixed distance $d_{push}$ to push the object. We assume there are $K$ candidate robot actions, i.e., $\mathcal{U}=\{\theta_{k}\}_{k=1}^{K}$. The value of each action is obtained through $\theta_{k}=2k\pi/K$ such that the starting positions of these $K$ candidate actions will evenly surround $\mathcal{S}_{\mathcal{A}}^{t}$.
As defined in Sec. 4, we require the object’s state to be always confined inside the cage region $\mathcal{S}_{cage}^{t}$, i.e., the PSS of the object $\mathcal{Q}_{t}\subset\mathcal{S}_{cage}^{t}$. Then, for the object to follow a certain trajectory $\mathcal{T}=\{(x^{d}_{1},x^{d}_{1}),\cdots,(x^{d}_{t},y^{d}_{t}),\cdots,(x^{d}_{T},y^{d}_{T})\}$ represented by a sequence of waypoints (i.e., desired positions) in the object’s state space, we just need to move the center of $\mathcal{S}_{cage}^{t}$, which is $(x^{c}_{t},y^{c}_{t})$, along the same trajectory. As the object is always confined inside the moving $\mathcal{S}_{cage}^{t}$ for every time step $t=1,\cdots,T$, the object’s position is guaranteed to follow the trajectory $\mathcal{T}$ with a positional error no greater than the cage size $R-r$.
@@ 5.2 Unknown Object Shape and Its Motion
For planar pushing, we represent the object’s motion by the displacement of the object’s position between adjacent time steps, denoted as $\mathbf{v}_{t}=\Delta\mathbf{q}_{t}=(\Delta x_{t},\Delta y_{t})$. Then the propagation function $\pi$, initially defined in Sec. 4.1, for a single object’s configuration and motion as inputs becomes $\pi(\mathbf{q}_{t},\mathbf{v}_{t})=\pi(\mathbf{q}_{t},\Delta\mathbf{q}_{t})=\mathbf{q}_{t}+\Delta\mathbf{q}_{t}$, which is used to propagate the object’s configuration $\mathbf{q}_{t+1}$ at the next step.
For a specific object’s configuration $\mathbf{q}_{t}$, the set of all possible motions $\mathcal{V}_{\mathbf{q}_{t}}$ is modeled by the interaction between the object and the robot’s line pusher $\mathcal{L}$ while the action $u_{t}$ is being executed, as represented by the function $U$ defined in Eq. (6). The pusher $\mathcal{L}$ approaches the object with a straight-line distance $d_{push}$ while $u_{t}$ is being executed. If the distance between $\mathbf{q}_{t}$ and the line pusher $\mathbf{d}(\mathbf{q}_{t},\mathcal{L})$ is greater than $r+d_{push}$, the object’s bounding circle will not collide with $\mathcal{L}$. The object will not move at all, and thus $\mathcal{V}_{\mathbf{q}_{t}}=\left\{(0,0)\right\}$. However, if the object’s bounding circle collides with $\mathcal{L}$, indicating the object is likely displaced by $\mathcal{L}$, we denote $d_{con}\leq d_{push}$ as the moving distance of $\mathcal{L}$ after contacting the bounding circle. As shown in Fig. 8, we can derive a bound for this displacement using the friction theory and Peshkin Distance (Peshkin and Sanderson, 1988) as follows:
$\mathcal{V}_{\mathbf{q}_{t}}=U(\mathbf{q}_{t},u_{t})=\text{if $\mathbf{d}(\mathbf{q}_{t},\mathcal{L})>r+d_{push}$ or $u_{t}$ is {None}:}$ (9)
@@
With this discretized representation, the PSS of the object is implemented by a finite set of pixels on $\mathcal{I}_{t}$. We also do a similar discretization for the possible motion set $\mathcal{V}_{\mathbf{q}_{t}}$. As such, to propagate the PSS on the discretized space, we just enumerate and propagate every pixel in the PSS with every motion in the discretized finite set of $\mathcal{V}_{\mathbf{q}_{t}}$. As mentioned, $\mathcal{I}_{t}$ is always centered with the moving cage. So we need to deal with an offset of $(x^{c}_{t},y^{c}_{t})$ whenever we convert between a pixel location on $\mathcal{I}_{t}$ and its actual coordinates in the original continuous space. Alg. 3 outlines the discretized implementation of the PSS propagation.
1
@@ 5.3 Cage the Push in Time
Recall that in Sec. 5.1, each candidate action in $\mathcal{U}$ is represented by an angle $\theta_{k}$, which determines the starting pose of the line pusher. For each candidate action in $\mathcal{U}$ with an index $k=1,\cdots,K$, we will evaluate a heuristic score by $h^{k}=\lambda_{1}S_{out}^{k}+\lambda_{2}\left(d_{out}^{k}\right)^{2}$ where $\lambda_{1}$ and $\lambda_{2}$ are weighting factors. As shown on the right of Fig. 7, $S_{out}^{k}$ is the POA area behind the line pusher when we place the pusher at a starting pose determined by the angle $\theta_{k}$, and $d_{out}^{k}$ is the furthest distance from behind the pusher to the POA boundary. It is worth noting that calculating the heuristics $h^{k}$ for each action does not require the propagation of PSS, i.e., the PSS remains $\mathcal{Q}_{t}$. Intuitively, a large $h^{k}$ indicates that a large portion of the PSS will go outside the cage if the $k$-th action is not taken.
@@ 6 Dynamic Tasks
In this section, we extend our Caging in Time theory from quasi-static tasks to dynamic scenarios where we instantiate the same framework on a challenging dynamic ball balancing problem. Tools are developed for propagating the ball’s PSS of its position and velocity, as well as for generating open-loop robot actions to keep the ball balanced while the robot end-effector follows different trajectories without requiring any sensory feedback. This is an instantiation of Caging in Time with an energy-based cage.
@@ 6.1 Problem Statement
While our framework employs a pure rolling assumption, this approach leverages a physical insight: rolling friction coefficients are typically smaller than sliding, introducing higher dynamic uncertainty. Our Caging in Time, specifically designed based on PSS, capitalizes on this property by effectively handling the more uncertain rolling dynamics, inherently encompassing all possible motion states, including sliding behaviors and transitional contact modes, without requiring explicit modeling of each interaction type.
@@
$E(\mathbf{q}_{t},\mathbf{g}_{\theta},\mathbf{a}_{p})=E_{k}(\mathbf{\dot{x}}_{t})+E_{ve}(\mathbf{x}_{t})+E_{p}(\mathbf{x}_{t},\mathbf{g}_{\theta},\mathbf{a}_{p})$ (12)
@@
$E_{max}(\mathbf{g}_{\theta},\mathbf{a}_{p})=E_{ve}(l)+E_{p}(l,\mathbf{g}_{\theta},\mathbf{a}_{p})$ (13)
@@ 6.2 Ball Dynamics and PSS Propagation
$=\int_{\mathbf{q}_{t}\in\mathcal{Q}_{t}}p(\mathbf{q}_{t})\int_{\mathbf{v}_{t}\in\mathcal{V}_{\mathbf{q}_{t}}}$
$\quad\quad\quad p(\mathbf{v}_{t}|\mathbf{q}_{t})\mathbbm{1}(\pi(\mathbf{q}_{t},\mathbf{v}_{t})=\mathbf{q}_{t+1})d\mathbf{v}_{t}d\mathbf{q}_{t}$
@@ 6.3 Cage the Ball in Time
To achieve Caging in Time for the dynamic ball balancing, we employ a control optimization strategy that combines Control Barrier Function (Ames et al., 2019) and Control Lyapunov Function (Anand et al., 2021) to keep the ball on the plate while following the desired trajectory.
In our case, Control Barrier Function (CBF) is utilized to implement the concept of “caging” in dynamic setting. As a physical cage constraining an object within a defined space, the CBF establishes an energy-based cage that can be characterized by the margin between the maximum energy attainable within the PSS and the preset energy boundary:
@@
Figure 14: Experiment setup. Left: Robot and camera setup. The robot performs the experiment on a transparent surface (blue) with a pre-installed line pusher (green). We used cameras 2 and 3 below the surface to record (not to track) the trajectories of the objects through April-Tag. Camera 1 was used to record the experiment from an upper view. Top Right: The 3D-printed objects used in the experiments. 1. Square with two trimmed tips, 2. Pentagon, 3. Triangle with trimmed tips, 4. Octagon, 5. Square, 6. Heart, 7. Hexagon, 8. Ellipse, 9. Circle, and letter-shaped objects: R, I, C, E. Bottom Right: 3D printed plates, tennis ball and plastic fruits for dynamic tasks. The square plate is for ball balancing when dimension $n=2$, and the thinner one is for ball balancing and catching when dimension $n=1$, which has a guide rail to constrain the motion of the ball only on the required dimension. The plastic fruits are strawberry (YCB #12), plum (YCB #18), pear (YCB #16) and lemon (YCB #14).